A population has a mean of 200 and a standard deviation of 50. Suppose a simple random sample of size 100 is selected and is used to estimate μ .
1. What is the probability that the sample mean will be within +/- 5 of the population mean (to 4 decimals)?

2. What is the probability that the sample mean will be within +/- 10 of the population mean (to 4 decimals)?

Answer by stanbon(75887)

A population has a mean of 200 and a standard deviation of 50.
Suppose a simple random sample of size 100 is selected and is
used to estimate μ .
-------------------------------
1. What is the probability that the sample mean will be within +/- 5 of the population mean (to 4 decimals)?
---
z(195) = (195-200)/50 = -0.1
z(205) = (205-200)/50 = +0.1
P(195 < x < 205) = p(-0.1 < z < 0.1) = 0.0797
-----------------------------------------------------
2. What is the probability that the sample mean will be within +/- 10 of the population mean (to 4 decimals)?
Do the same using 190 and 210
Ans: 0.1585
==================
Cheers,
Stan H.
==================
-----------------------------------